# Solving Empirical Bayes via Transformers

### v1

@@ Abstract
This work applies modern AI tools (transformers) to solving one of the oldest statistical problems: Poisson means under empirical Bayes (Poisson-EB) setting. In Poisson-EB a high-dimensional mean vector $\theta$ (with iid coordinates sampled from an unknown prior $\pi$) is estimated on the basis of $X=\mathrm{Poisson}(\theta)$. A transformer model is pre-trained on a set of synthetically generated pairs $(X,\theta)$ and learns to do in-context learning (ICL) by adapting to unknown $\pi$. Theoretically, we show that a sufficiently wide transformer can achieve vanishing regret with respect to an oracle estimator who knows $\pi$ as dimension grows to infinity. Practically, we discover that already very small models (100k parameters) are able to outperform the best classical algorithm (non-parametric maximum likelihood, or NPMLE) both in runtime and validation loss, which we compute on out-of-distribution synthetic data as well as real-world datasets (NHL hockey, MLB baseball, BookCorpusOpen). Finally, by using linear probes, we confirm that the transformer’s EB estimator appears to internally work differently from either NPMLE or Robbins’ estimators.
@@ 1 Introduction
Transformers have received a lot of attention due to the prevalence of large language models (LLM). More generally, we think of (encoder-only) transformers as generic engines for learning from exchangeable data. Since most classical statistical tasks are formulated under iid sampling assumption, it is very natural to try to apply transformers to them [GTLV22].
Training transformers for classical statistical problems serves two purposes. One is obviously to get better estimators. Another, equally important, goal of such exercises is to elucidate the internal workings of transformers in a domain with a much easier and much better understood statistical structure than NLP. In this work, we believe, we found the simplest possible such statistical task: empirical Bayes (EB) mean estimation. We believe transformers are suitable for EB because EB estimators naturally exhibit a shrinkage effect (i.e. biasing mean estimates towards the nearest mode of the prior), and so do transformers, as shown in [GLPR24] that the attention mechanisms tend to cluster tokens. Additionally, the EB mean estimation problem is permutation equivariant, removing the need for positional encoding. In turn, estimators for this problem are in high demand [KG24, GK23, GK22] and unfortunately the best classical estimator (so-called non-parametric maximum likelihood, or NPMLE) suffers from slow convergence. In this work, we demonstrate that transformers outperform NPMLE while also running almost 100x faster. We now proceed to defining the EB task.
Poisson-EB task: One observes $n$ samples $X_{1},\ldots,X_{n}$ which are generated iid via a two-step process. First, $\theta_{1},\ldots,\theta_{n}$ are sampled from some unknown prior $\pi$ on $\mathbb{R}$. The $\pi$ serves as an unseen (non-parametric) latent variable and we assume nothing about it (not even continuity or smoothness). Second, given $\theta_{i}$’s, we sample $X_{i}$’s conditionally iid via $X_{i}\sim\text{Poi}(\theta_{i})$. The goal is to estimate $\theta_{1},\cdots,\theta_{n}$ via $\hat{\theta}_{1},\cdots,\hat{\theta}_{n}$ upon seeing $X_{1},\cdots,X_{n}$ that minimizes the expected mean-squared error (MSE), $\mathbb{E}[(\hat{\theta}(X)-\theta)^{2}]$. If $\pi$ were known, the Bayes estimator that minimizes the MSE is the posterior mean of $\theta$, which also has the following form.
@@
$\mathsf{Regret}(\hat{\theta})$ $=\mathbb{E}\left[\left(\hat{\theta}(X)-\theta\right)^{2}\right]-\mathbb{E}\left[\left(\theta_{\pi}(X)-\theta\right)^{2}\right]$
$=\mathbb{E}\left[\left(\hat{\theta}(X)-\theta\right)^{2}\right]-\mathsf{mmse}(\pi)$
In this Poisson-EB setting, multiple lines of work have produced estimators that resulted in regret that vanishes as sample size increases [BGR13, PW21, JPW22, JPTW23]. Robbins estimator [Rob51, Rob56] replaces the unknown posterior density $f_{\pi}$ in (1) with $N_{n}(\cdot)$, the empirical count among the samples $X_{1},\cdots,X_{n}$. Minimum distance estimators first estimate a prior (e.g. the NPMLE estimator $\hat{\pi}_{\mathsf{NPMLE}}=\mathop{\rm argmax}_{Q}\prod_{i=1}^{n}f_{Q}(X_{i})$), and then produces the plugged-in Bayes estimator $\hat{\theta}_{\hat{\pi}}$. Notice that Robbins estimator suffers from multiple shortcomings like numerical instability (c.f. [EH21]) and the lack of monotonicity property of Bayes estimator $\hat{\theta}_{\pi}$ (c.f. [HS83]), while minimum-distance estimators are too computationally expensive and do not scale to higher dimensions. [JPTW23] attempts to remedy the ‘regularity vs efficiency’ tradeoffs in these estimators with an estimator based on score-estimation equivalent in the Poisson model. However, despite the monotone regularity added, this estimator still does not have a Bayesian form: a cost one pays to achieve an efficient computational time.
Solving Poisson-EB via transformers. We formulate our procedure for solving Poisson-EB as follows: we generate synthetic data and train our transformers on those. Then, we freeze their weights and present new data to be estimated. To our knowledge, this is the first line of work that studies using neural network models for empirical Bayes. Concretely, our contributions are as follows:
@@
    In Section 4, we show that transformers can approximate Robbins and the NPMLE via the universal approximation theorem. We also use linear probes to show that our pre-trained transformers work differently than the two aforementioned estimators.
@@
    In Section 5, we set up synthetic experiments to demonstrate that synthetically pre-trained transformers can generalize to unseen sequence lengths and evaluation priors. This is akin to [XRLM21] where ICL occurs at test time despite distribution mismatch.
@@
One significance of our synthetic experiments is that transformers demonstrate length-generalization by achieving lower regret upon being tested on sequence lengths up to 4x the length they are trained on, even on unseen priors. This comes as multiple works show mixed success of length-generalization of transformers [ZAC+24, WJW+24, KPNR+24, AWA+22].
We mention that there is a long literature studying transformers for many statistical problems [BCW+24]. What makes this work different is that a) our estimator does not just match but improve upon existing (classical) estimators, thus advancing the statistical frontier; b) our setting is unsupervised and much closer to NLP compared to most previous work considering supervised learning (classification and regression), in which data comes in pairs, thus requiring unnatural tricks to pair tokens; c) our problem is non-parametric.
In summary, we demonstrate that even for classical statistical problems, transformers offer an excellent alternative (in runtime and performance). For the simple 1D Poisson-EB task, we also found that already very economically sized transformers ($<100$ k parameters) can have excellent performance.
@@ 2 Related Work
Transformers and in-context learning (ICL). Transformers have shown the ability to do ICL, as per the thread of work summarized in [DLD+22]. ICL is primarily manifested in natural language processing [BMR+20, DSD+22] and learning linear models [ASA+22, ZFB23]. Other examples that transformers can learn are gradient descent [BCW+24], several non-linear function classes [GTLV22], and support vector machine [TLTO23], while having limited ability on boolean functions [BPBK23]. Recent works have also explained ICL from the Bayesian point of view [MHH24, PAG23], including showing Bayesian behavior even upon train-test distribution mismatch [XRLM21].
@@
How do transformers work? [YBR+19] have established the universal approximation theorem of transformers. This was later extended for sparse transformers [YCB+20] and ICL setting [FdHP24]. Its limitations are further discussed in [NKB24]. Transformers have also been shown to do other approximation tasks, like Turing machines [WCM22, PBM21]. From another perspective, [AB18] introduces linear probes as a mechanism of understanding the internals of a neural network, which is further studied in [Bel22]. Linear probe has also been applied in transformers to study its ability to perform NLP tasks [TDP19], achieve second order convergence [FCJS24], and learn various functions in-context [GHM+23]. One such application is ICL linear regression to look for moments [ASA+22]. Recently, linear probe has been used by [AZR+24] to improve in-context learning.
Empirical Bayes. Empirical Bayes is a powerful tool for large-scale inference [Efr12]. Some of its applications include performing downstream tasks like linear regression [KWCS24, MSS23], estimating the number of missing species [FCW43], and large scale hypothesis testing [ETST01]. In computational biology, empirical Bayes has also been used in sequencing frameworks [HK10, LDT+13], though these frameworks are mostly parametric and rely on estimating the parameters of a prior.
In the theoretical setting, multiple lines of work have established the theoretical bounds that can be achieved by empirical Bayes estimators. In the Poisson-EB problem, Robbins [Rob51, Rob56] formulated an estimator based on Tweedie’s formula, known as $f$-modelling. In the normal means EB problem, [JZ09] formulated a $g$-modelling approach via prior estimation, which was also adapted to the Poisson-EB problem. More recently, [JPTW23] formulated an estimator based on ERM on monotone functions, which introduces regularity to the estimators while also escaping the computationally expensive prior estimation process. The optimality of these estimators has been established in the following works: [BGR13, PW21, JPW22, JPTW23].
@@ 3.1 Baselines description
Non empirical Bayes baselines. When nothing is known about the prior $\pi$ the minimax optimal estimator is the familiar maximum-likelihood (MLE) estimator $\hat{\theta}_{\mathsf{MLE}}(x)=x$. However, when one restricts priors in some way, the minimax optimal estimator is not MLE, but rather a Bayes estimator for the worst-case prior. In this work, we consider priors restricted to support $[0,50]$. The minimax optimal estimator for this case is referred to as the gold standard (GS) estimator to signify its role as the “best” in the sense of classical (pre-EB) statistics. Appendix A.1 contains derivation of GS.
Empirical Bayes baselines. We will use the following empirical Bayes estimators as introduced in Section 1: the Robbins estimator, NPMLE estimator, and the ERM-monotone estimator with algorithm described in Lemma 1 of [JPTW23].
@@ 3.2 Transformer Architecture
Next, we describe our transformer architecture, which closely mimics the standard transformer architecture in [VSP+17]. Given the permutation invariance of the Bayes estimator, we do not use positional encoding or masking. Thus effectively, it is a full-attention encoder-only transformer with one linear decoder on top.
One aspect worth mentioning is that at the encoding stage, we are using two different weights, split evenly across the $N$ layers. The intuition behind it is that one learns the encoding part (input) and the other the decoding part (output).
@@ 3.3 Training Protocol
Data generation. We emphasize that all our transformers are trained on synthetic data, using the Poisson-generated integers $X$ as inputs and the hidden parameters $\theta$ as labels. We use the plain vanilla MSE loss $\sum(\hat{\theta}(X_{i})-\theta_{i})^{2}$. There are two classes of priors from which we generate $\theta$, the neural-generated prior-on-priors, and Dirichlet process with base distribution $\mathsf{Unif}[0,50]$ within each batch. We fix the sentence length = 512 throughout training. With the exception as noted later in Section 5.1, we cap the label at $\theta_{\max}=50$ (i.e. our priors are in the class $\mathcal{P}([0,50])$. We defer the detailed discussion to Appendix A.2, including the motivation to train with a mixture of the two priors.
Parameter Selection. We consider models of 6, 12, 18, 24, and 48 layers, embedding dimension $\mathsf{dmodel}$ either 32 or 64, and number of heads in 4, 8, 16, 32. We fix the number of training epochs to 50k, the learning rate to 0.02, and the decay rate every 300 epochs to 0.9. Among the trained models, we chose our models based on the mean-squared error evaluated on neural prior-on-prior and Dirichlet process during inference time. We then arrive at the two models described in Table 1, which we will name T18 and T24 depending on their number of layers. Both have around 25.6k parameters. We also define T18r and T24r as the transformers we train with random $\theta_{\max}$.  [^1] In the future, we will add T18r and T24r to all comparisons, but for now they only appear on Fig. 5 and Section 5.1.
Table 1: The characteristics of T18 and T24, respectively.
Transformer Layers Embedding dimension # Heads $\theta_{\max}$
T18 18 32 16 50
T24 24 32 8 50
T18r 18 32 16 Random $[10,150]$
T24r 24 32 8 Random $[10,150]$

### v2

@@ Abstract
This work applies modern AI tools (transformers) to solving one of the oldest statistical problems: Poisson means under empirical Bayes (Poisson-EB) setting. In Poisson-EB, a high-dimensional mean vector $\theta$ (with i.i.d. coordinates sampled from an unknown prior $\pi$) is estimated on the basis of $X=\mathrm{Poisson}(\theta)$. A transformer model is pre-trained on a set of synthetically generated pairs $(X,\theta)$ and learns to do in-context learning (ICL) by adapting to unknown $\pi$. Theoretically, we show that a sufficiently wide transformer can achieve vanishing regret with respect to an oracle estimator that knows $\pi$ as the dimension grows to infinity. Practically, we discover that already very small models (100k parameters) are able to outperform the best classical algorithm (non-parametric maximum likelihood, or NPMLE) both in runtime and validation loss, which we compute on out-of-distribution synthetic data as well as real-world datasets (NHL hockey, MLB baseball, BookCorpusOpen). Finally, by using linear probes, we confirm that the transformer’s EB estimator appears to internally work differently from either NPMLE or Robbins’ estimators.
@@ 1 Introduction
Transformers have received a lot of attention due to the prevalence of large language models (LLM). More generally, we think of (encoder-only) transformers as generic engines for learning from exchangeable data. Since most classical statistical tasks are formulated under the i.i.d. sampling assumption, it is very natural to try to apply transformers to them [GTLV22].
Training transformers for classical statistical problems serves two purposes. One is obviously to get better estimators. Another, equally important, goal of such exercises is to elucidate the internal workings of transformers in a domain with a much easier and much better understood statistical structure than NLP. In this work, we believe, we found the simplest possible such statistical task: empirical Bayes (EB) mean estimation. We believe transformers are suitable for EB because EB estimators naturally exhibit a shrinkage effect (i.e. biasing mean estimates towards the nearest mode of the prior), and so do transformers, as shown in [GLPR24] that the attention mechanisms tend to cluster tokens. Additionally, the EB mean estimation problem is permutation equivariant, removing the need for positional encoding. In turn, estimators for this problem are in high demand [KG24, GK23, GK22] and unfortunately, the best classical estimator (so-called non-parametric maximum likelihood, or NPMLE) suffers from slow convergence. In this work, we demonstrate that transformers outperform NPMLE while also running almost 100x faster. We now proceed to defining the EB task.
Poisson-EB task: One observes $n$ samples $X_{1},\ldots,X_{n}$ which are generated iid via a two-step process. First, $\theta_{1},\ldots,\theta_{n}$ are sampled from some unknown prior $\pi$ on $\mathbb{R}$. The $\pi$ serves as an unseen (non-parametric) latent variable, and we assume nothing about it (not even continuity or smoothness). Second, given $\theta_{i}$’s, we sample $X_{i}$’s conditionally iid via $X_{i}\sim\text{Poi}(\theta_{i})$. The goal is to estimate $\theta_{1},\cdots,\theta_{n}$ via $\hat{\theta}_{1},\cdots,\hat{\theta}_{n}$ upon seeing $X_{1},\cdots,X_{n}$ that minimizes the expected mean-squared error (MSE), $\mathbb{E}[(\hat{\theta}(X)-\theta)^{2}]$. If $\pi$ were known, the Bayes estimator that minimizes the MSE is the posterior mean of $\theta$, which also has the following form.
@@
$\mathsf{Regret}(\hat{\theta})=\mathbb{E}\left[\left(\hat{\theta}(X)-\theta\right)^{2}\right]-\mathbb{E}\left[\left(\theta_{\pi}(X)-\theta\right)^{2}\right]=\mathbb{E}\left[\left(\hat{\theta}(X)-\theta\right)^{2}\right]-\mathsf{mmse}(\pi)$
In this Poisson-EB setting, multiple lines of work have produced estimators that resulted in regret that vanishes as sample size increases [BGR13, PW21, JPW22, JPTW23]. Robbins estimator [Rob51, Rob56] replaces the unknown posterior density $f_{\pi}$ in (1) with $N_{n}(\cdot)$, the empirical count among the samples $X_{1},\cdots,X_{n}$. Minimum distance estimators first estimate a prior (e.g. the NPMLE estimator $\hat{\pi}_{\mathsf{NPMLE}}=\mathop{\rm argmax}_{Q}\prod_{i=1}^{n}f_{Q}(X_{i})$), and then produces the plugged-in Bayes estimator $\hat{\theta}_{\hat{\pi}}$. Notice that Robbins estimator suffers from multiple shortcomings like numerical instability (c.f. [EH21]) and the lack of monotonicity property of the Bayes estimator $\hat{\theta}_{\pi}$ (c.f. [HS83]), while minimum-distance estimators are too computationally expensive and do not scale to higher dimensions. [JPTW23] attempts to remedy the ‘regularity vs efficiency’ tradeoffs in these estimators with an estimator based on score-estimation equivalent in the Poisson model. However, despite the monotone regularity added, this estimator still does not have a Bayesian form: a cost one pays to achieve an efficient computational time.
Solving Poisson-EB via transformers. We formulate our procedure for solving Poisson-EB as follows: we generate synthetic data and train our transformers on them. Then, we freeze their weights and present new data to be estimated. Concretely, our contributions are as follows:
@@
    In Section 4, we show that transformers can approximate Robbins and the NPMLE via the universal approximation theorem. We also use linear probes to show that our pre-trained transformers appear to work differently from the two aforementioned classical EB estimators.
@@
    In Section 5, we set up synthetic experiments to demonstrate that pre-trained transformers can generalize to unseen sequence lengths and evaluation priors. This is akin to [XRLM21] where ICL occurs at test time despite distribution mismatch. In addition, we will show that a transformer of reduced complexity can perform at almost the same level.
@@
One interesting finding is that our transformers demonstrate length-generalization, whereas previously transformers’ track record in this regard was rather mixed [ZAC+24, WJW+24, KPNR+24, AWA+22]. Specifically, our transformers achieve length-generalization by getting lower regret upon being tested on sequence lengths up to 4x the length they were trained on, even in zero-shot setting where the data comes from unseen type of prior.
We mention that there is a long literature studying transformers for many statistical problems [BCW+24], further reviewed below. What distinguishes our work is a) our estimator does not just match but improve upon existing (classical) estimators, thus advancing the statistical frontier; b) our setting is unsupervised and closer to NLP compared to most previous work considering supervised learning (classification and regression), in which data comes in pairs, thus requiring unnatural tricks to pair tokens; c) our problem is non-parametric.
In summary, we demonstrate that even for classical statistical problems, transformers offer an excellent alternative (in runtime and performance). For the simple 1D Poisson-EB task, we also found that already very economically sized transformers ($<100$ k parameters) can have excellent performance. We hope that disecting the inner workings of these transformers will suggest novel types of statistical procedures for the large-scale statistical inference [Efr12].
@@ 2 Related work
Transformers and in-context learning (ICL). Transformers have shown the ability to do ICL, as per the thread of work summarized in [DLD+22]. ICL is primarily manifested in natural language processing [BMR+20, DSD+22]; several works also study the ability and limitations of transformers in using ICL in doing various statistical tasks [ASA+22, ZFB23, GTLV22, TLTO23, BPBK23, ASA+22, BCW+24]. Recent works have also explained ICL from a Bayesian point of view [MHH24, PAG23], including upon train-test distribution mismatch [XRLM21].
@@
How do transformers work? [YBR+19] has established the universal approximation theorem of transformers. This was later extended for sparse transformers [YCB+20] and the ICL setting [FdHP24]. Its limitations are further discussed in [NKB24]. Transformers have also been shown to do other approximation tasks, like Turing machines [WCM22, PBM21]. From another perspective, [AB18] introduces linear probes as a mechanism of understanding the internals of a neural network. Linear probe has also been applied in transformers to study its ability to perform NLP tasks [TDP19], achieve second order convergence [FCJS24], learn various functions in-context [GHM+23], and improve in-context learning [AZR+24].
Empirical Bayes. Empirical Bayes was introduced in [Rob56] based on the premise that estimation tasks on a sequence can yield lower risk by allowing estimation of each individual component to depend on the entire sequence, e.g. as demonstrated by the James-Stein estimator [Ste56, JS+61]. In the Poisson model, known estimators for the mean estimation problem are based either on the Tweedie formula [Rob56], posterior density estimation [KW56, Lin83], or ERM [JPTW23]. These methods have been shown to yield minimax optimal regret for certain classes of priors [BGR13, PW21, JPW22, JPTW23]. In practice, empirical Bayes is a powerful tool for large-scale inference [Efr12], with applications including downstream tasks like linear regression [KWCS24, MSS23], estimating the number of missing species [FCW43], large scale hypothesis testing [ETST01], and the construction of biological sequencing frameworks [HK10, LDT+13].
Comparison to Bayesian methods and SBI. We compare our approach against simulation-based inference (SBI) [CBL20] in the fully Bayesian model. In both EB and Bayesian model, we desire to obtain posterior $p(\theta|X)$, which requires knowledge of a good prior $\pi$ on $\theta$. In SBI, this is resolved via a method of amortized inference [ZMSDH24], where one trains a model on many $(\theta,x)$ pairs thus implicitly learning the prior (and often the forward model $p(X|\theta)$ as well). It is hoped that the learned prior matches the one during inference. We note that transformers have been applied in Bayesian settings [HMP+25, GDW+24, CLH+25], but these applications require either labeled (supervised) data at inference time or simulation access to the prior $\pi$ itself. Unsurprisingly, prior misspecification causes serious problems [ZMSDH24, Appendix S2.8] unless one uses some regularization [HBS+23].
In contrast, in the EB setting, absolutely nothing about the prior is assumed at inference time, with the model instead trying to estimate $\mathbb{E}[\theta|X=x]$ directly from the data consisting of $X_{i}$’s only. We stress again, that in EB there are no labeled (no $\theta$’s!) data presented to the estimator at inference time. Thus, while our model is trained on synthetic examples sampled from a variety of priors, the prior at inference time is expected to deviate from those used at pre-training and the whole goal of our work is to improve the zero-shot generalization ability. We will discuss more in Appendix F.1, using illustrations of real-life datasets where this comparison becomes more apparent.
@@ 3.1 Baselines description
Non empirical Bayes baselines. When nothing is known about the prior $\pi$, the minimax optimal estimator is the familiar maximum-likelihood (MLE) estimator $\hat{\theta}_{\mathsf{MLE}}(x)=x$. However, when one restricts priors in some way, the minimax optimal estimator is not MLE, but rather a Bayes estimator for the worst-case prior. In this work, we consider priors restricted to support $[0,50]$. The minimax optimal estimator for this case is referred to as the gold standard (GS) estimator to signify its role as the “best” in the sense of classical (pre-EB) statistics. Appendix A.1 contains derivation of GS.
Empirical Bayes baselines. We will use the following empirical Bayes estimators as introduced in Section 1: the Robbins estimator, NPMLE estimator, and the ERM-monotone estimator with the algorithm described in Lemma 1 of [JPTW23].
@@ 3.2 Transformer architecture
Next, we describe our transformer architecture, which closely mimics the standard transformer architecture in [VSP+17]. Given the permutation invariance of the Bayes estimator, we do not use positional encoding or masking. Thus effectively, it is a full-attention encoder-only transformer with one linear decoder on top. In addition, we achieve parameter sharing by using two different weights, split evenly across the $N$ layers. The intuition behind it is that one learns the encoding part (input) and the other the decoding part (output).
Linear attention. We considered the following linearized attention using some feature representation $\phi$:
$\text{Attention}(Q,K,V)=\frac{1}{n}\phi(Q)(\phi(K)^{T}V)$ (2)
We experimented with several choices of $\phi$ and found that the identity function $\phi(x)=x$ works well.
Linearized attention is formalized in [KVPF20], and subsequently refined in frameworks like Mamba [GD23, DG24], gated update rule [YWS+23], and Delta rule [YKH24, YWZ+24]. While we will show that the attention defined in (2) performs almost as well in our EB setting, a rigorous comparison of different linear attention frameworks in solving EB will be subject to future work.
@@ 3.3 Training protocol
Data generation. We emphasize that all our transformers are trained on synthetic data, using the Poisson-generated integers $X$ as inputs and the hidden parameters $\theta$ as labels. We train our transformers via the MSE loss $\sum_{i=1}^{n}(\hat{\theta}(X_{i})-\theta_{i})^{2}$ and the Adam optimizer [Kin14].
There are two classes of priors from which we generate $\theta_{\text{base}}\in[0,1]$: the neural-generated prior-on-priors $\in\mathcal{P}([0,1])$, and Dirichlet process with base distribution $\mathsf{Unif}[0,1]$ within each batch. These are further described in Appendix A.2. We then define $\theta=\theta_{\text{base}}\cdot\theta_{\max}$. At evaluation stage, $\theta_{\max}$ is generally fixed (e.g. 50 in Section 5). To ensure robustness of our transformers against inputs of different magnitude, however, at training stage we let $\theta_{\max}$ be random, sampled from
$\frac{3}{4}\text{Unif}([0,200])+\frac{1}{8}\text{Exp}(50)+\frac{1}{8}\text{Cauchy}(50,10)$
and capped at $\theta_{\max}\leq 500$. We defer the detailed discussion to Appendix A.2, including the motivation to train with a mixture of the two priors.
Parameter selection. We consider models of 6, 12, 18, 24, and 48 layers, embedding dimension $\mathsf{dmodel}$ either 32 or 64, and number of heads in 4, 8, 16, 32. We fix the number of training epochs to 50k, the learning rate to 0.02, and the decay rate every 300 epochs to 0.9. We train our transformers on sequence length 512, and for each epoch the we use 192 batches. Among the trained models, we chose our models based on the mean-squared error evaluated on neural prior-on-prior and Dirichlet process during inference time. We eventually arrive at our model T24r, with 24 layers, embedding dimension 32, 8 heads. We will make comparison against L24r, which is its linear attention counterpart as defined in (2), and with the identity feature representation $\phi(x)=x$.